What is the prime factorization of 385 Please show the steps
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

What is the prime factorization of 385 Please show the steps
-------------
385 = 5*77 = 5*7*11
=========================
Cheers,
Stan H.

Answer by richwmiller(17219)

we know that 5 is a factor
so let's pull out 5
385/5=77
I think that's obvious
7 and 11
and we are done
5,7,11